OpenStudy (anonymous):

find the remainder when f(x)=2x^3-12x^2+11x+2 is divided by x-5?

OpenStudy (anonymous):

find f(5)

OpenStudy (anonymous):

i got 7

OpenStudy (anonymous):

if you did the math right, then 7 it is

OpenStudy (anonymous):

can u tell me if im right or wrong so if not i can try again?

OpenStudy (akashdeepdeb):

There are 2 ways in which you can solve this. # The remainder theorem method. # The long division method. This is what the first method says: If f(x) is divisible by (x-a), then f(a) = 0. Or when f(x) is divided by (x-a) then f(a) = b is the remainder. And vice versa. The second way is just normal long division of polynomials. The shorter and better method to solve this would be the remainder theorem method. So is f(x)=2x^3-12x^2+11x+2 is divided by x-5? So let us see what the remainder is: f(x) = 2x^3-12x^2+11x+2 f(5) = 2*5^3-12*5^2+11*5+2 f(5) = 250 - 300 + 55 + 2 = 7 Thus the remainder is 7 !! :D Understood? :)

OpenStudy (anonymous):

thnk you!!! yes i think ive got it :)

OpenStudy (akashdeepdeb):

:)